OpenStudy (anonymous):

Use implicit differentiation to find slope of tangent line to the curve xy-y^2+4x=-2 at the point (1,-2)

OpenStudy (irishboy123):

\(xy-y^2+4x=-2\) so: \(\frac{d}{dx}(xy)-\frac{d}{dx}(y^2)+\frac{d}{dx}(4x)=\frac{d}{dx}(-2)\) what do you get?